You are selling tickets for a high school concert. student tickets cost $4 and general admission tickets cost $6. You sell 450 tickets and collect $2340. How many of each type of ticket did you sell?
Answer by ptaylor(1894)


Using 2 unknowns:
Let x=number of student tickets sold
And let y=number of general admission tickets sold
Now we are told the following:
x+y=450----------------------------------eq1
and($ understood)
4x+6y=2340----------------------------eq2
multiply eq1 by 4:
4x+4y=1800------------------------------------eq1a
subtract eq1a from eq2 and we get:
2y=540 divide boh sides by 2
y=270--------------------------number of general admission tickets sold
Substitute y=270 into eq 1:
x+270=450 subtract 260 from each side
x+270-270=450-270 collect like terms
x=180------------------------------------number of student tickets sold
CK
180+270=450
450=450
and
4*180+6*270=2340
720+1620=2340
2340=2340
Using 1 unknown
Let x=number of student tickets sold
Then 450-x=number of general admission tickets sold
Now we are told that:
($ understood)
4x+6(450-x)=2340 get rid of parens (distributive law)
4x+2700-6x=2340 subtract 2700 from each side
4x+2700-2700-6x=2340-2700 collect like terms
-2x=-360 divide both sides by -2
x=180-----------------------number of student tickets sold
450-x=450-180=270------------------number of general admission tickest sold

Hope this helps---ptaylor